OpenStudy (anonymous):

The half lie of randon-222 is 3.8 days. How much of a 100g sample is left after 15.2 days?

OpenStudy (agl202):

15.2 days = 15.2/3.8 = 4 half days Then leftover = 100 grams * (0.5)^4 = 6.25 grams

OpenStudy (isaiah.feynman):

Sorry, I'm late!